OpenStudy (anonymous):

Find the 13th term of the following sequence -3,-18,-108... a: 6*(-3)^13 c: -3*6^13 b: 6*(-3)^12 d: -3*6^12

OpenStudy (anonymous):

it is a G.P. a=-3 \[r=\frac{ -18 }{ -3 }=6\] \[t _{n}=ar ^{n-1}\] put n=13 and get the solution.

OpenStudy (anonymous):

So its -3*6^12?

OpenStudy (texaschic101):

tn = ar^(n-1) t13 = -3* 6^(13 - 1) = t13 = -3 * 6^12 yep..you got it